K= 2b + 5c; solve b then solve c. I am very rusty in algebra. My son and I are getting different answers and he doesn't have his book so I can refresh my memory. Can you work this out so we can get and example to finish the others? Thank you.
Answer by edjones(8007)

k=2b+5c
.
solve for b.
k-2b=5c
-2b=5c-k
b=(5c-k)/-2
b=-(5c-k)/2
b=(k-5c)/2
.
solve for c.
k-5c=2b
-5c=2b-k
c=(2b-k)/-5
c=-(2b-k)/5
c=(k-2b)/5
.
Ed